OpenStudy (anonymous):

Mr. Simon landscapes new houses for a building contractor. He charges by the square meter of yard space plus a setup fee. He billed the contractor $2,150 for a 400-square- meter yard and $1,650 for a 300-square-meter yard. Write an equation in the form y = mx + b that can be used to find the cost, y, to landscape any amount of yard space, x.

OpenStudy (dtan5457):

You can make a system to find y 400m+b=2150 300m+b=1650

OpenStudy (dtan5457):

Solve for m, to get b.

OpenStudy (dtan5457):

Good from here..?

OpenStudy (anonymous):

not really @dtan5457 im new to this subject i have it as an online class

OpenStudy (dtan5457):

400m+b=2150 300m+b=1650 subtract everything to get 100m=500 m=?

OpenStudy (dtan5457):

Still don't get it?

OpenStudy (dtan5457):

Divide everything by 100.

OpenStudy (anonymous):

ok @dtan5457

OpenStudy (dtan5457):

So m=?